Circular No. 7048

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J0111.2–7317 D. Chakrabarty, A. M. Levine, and G. W. Clark, Massachusetts In-

stitute of Technology; and T. Takeshima, Goddard Space Flight Center, NASA, and Universities Space Research Association, write: “Multiple observations of the SMC X-1 field with the Proportional Counter Array (PCA) on the Rossi X-ray Timing Explorer between Oct. 29 and Nov. 7 have revealed a previously unknown 31-s x-ray pulsar, designated XTE J0111.2– 7317, ∼ 30′ from SMC X-1. Raster scans of the region on Nov. 6 and 7 have yielded the following position for the new pulsar: α = 1^{h}11^{m}.2, δ = −73^{o}17′.4 (equinox 2000.0), with an error radius of 1′.5 (90-percent confidence). The 2–10-keV x-ray intensities of XTE J0111.2–7317 and SMC X-1 on Nov. 6.6 UT were 36(7) and 51(4) mCrab, respectively, and the pulsed fraction of XTE J0111.2–7317 was roughly 45 percent. If XTE J0111.2–7317 is located in the SMC, its x-ray intensity implies a luminosity of order 10^{38} erg/s. The source was not detected during a series of six PCA observations of the field between Oct. 16 and 27. Observations at other wavelengths are encouraged in order to identify the binary companion.”

C. A. Wilson and M. H. Finger, Marshall Space Flight Center, NASA, report: “Hard-x-ray observations with the BATSE all-sky monitor on the Compton Gamma Ray Observatory first detected XTE J0111.2–7317 on Oct. 30, with a barycentric pulse frequency of 0.0322275(8) Hz on Oct. 30.5 UT. The 20–50-keV pulsed intensity increased from 18(3) mCrab on Oct. 30 to 37(3) mCrab on Nov. 5 and remained steady through Nov. 9. The pulse profile in the band 20–50 keV consists of a single asymmetric peak. BATSE pulse-phase measurements during Nov. 2–8 are consistent with a barycentric pulse frequency of 0.03225181(3) Hz on Nov. 5.0 and a constant pulse frequency derivative of 5.58(4) × 10^{−11} Hz/s. If this rate is primarily due to the accretion torque on the pulsar, it implies a luminosity of at least 10^{38} erg/s, consistent with the distance to the SMC.”

COMET C/1998 U5 (LINEAR) Total visual magnitude and coma-diameter estimates: Nov. 7.76 UT,

10.0, 2′.8 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 8.72, 10.2, 1′.0 (T. Karhula, Virsbo, Sweden, 0.20-m reflector); 9.87, 10.2, 2′.5 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 10.89, 9.9, 3′.0 (Bouma).

1998 November 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green